OpenStudy (anonymous):

Is there a method where I'm able to solve \(7-r-2k=0\) without trial and error?

OpenStudy (unklerhaukus):

solve for r(k) or k(r)?

OpenStudy (anonymous):

solve for both \(r\) and \(k\)

OpenStudy (anonymous):

like finding both variable.

OpenStudy (unklerhaukus):

\[ 7−r−2k=0 \] \[ r=7−2k \]\[k=\frac{r-7}2\]

OpenStudy (anonymous):

Omniscience, this equation doesn't have a single answer. It has infinitely many answers, which is why we can't simply say, "Oh, r equals this, and k equals this." So instead, what we can do is exactly what my boy Rhaukus suggested. We can find an equation r(k) or k(r) which means then that we can pick any k, plug it into the equation to get r, and there would be ONE of the many solutions.

OpenStudy (unklerhaukus):

\[*\qquad k=\frac{7-r}2\]

OpenStudy (anonymous):

okay, then i can just the first variable and find the next?

OpenStudy (anonymous):

find the first*

OpenStudy (anonymous):

So, Rhaukus solved for r and got: r = 7-2k To find an individual solution, I'll pick any k that I want. I want to pick k=5, just because. I plug k into my equation: r = 7-2(5) r= 7-10 r=-3 So ONE solution of the equation is k=5, r=-3

OpenStudy (unklerhaukus):

the value of the variables depend on each other

OpenStudy (anonymous):

But its from this.. \[\left(2x+3\right)^{4} \left(x-\frac{2}{x}\right) ^{2} \] \[\left(\begin{matrix}4 \\ r\end{matrix}\right)*\left(\begin{matrix}3 \\ k\end{matrix}\right)2^{4-r}3^{r}(-2)^{k}x^{7-r-2k}\] where im finding the variable of \(r\) and \(k\)

OpenStudy (anonymous):

ahaha oh my god.

OpenStudy (anonymous):

which means im only able to use trial and error? because you would never know which number to plug in..

OpenStudy (unklerhaukus):

substitute \(r=7-2k\) into this

OpenStudy (unklerhaukus):

\[\left(\begin{matrix}4 \\ 7−2k \end{matrix}\right)*\left(\begin{matrix}3 \\ k\end{matrix}\right)2^{4-7−2k }3^{7−2k }(-2)^{k}x^{7-(7-2k)-2k}\]

OpenStudy (unklerhaukus):

well the coefficient of x can be simplified

OpenStudy (anonymous):

is that method allowed?

OpenStudy (unklerhaukus):

i dont know the context, im not even sure what is in those big brackets...

OpenStudy (anonymous):

well, im finding the constant term for the expansion..

OpenStudy (unklerhaukus):

does it depend on x?

OpenStudy (anonymous):

I am just wondering; if there is a method where I don't need to use trial and error because it's long. Thanks anyway!

OpenStudy (anonymous):

The constant term does not have \(x\) in it; the independent term